OpenStudy (anonymous):

find cos Θ if sin Θ= -2/3 and falls in quadrant 4

OpenStudy (ankit042):

sin^x+cos^x =1

OpenStudy (ankit042):

also cos is +ve in 4th quadrant

OpenStudy (anonymous):



OpenStudy (anonymous):

find the third side of the triangle via pythagoras in your head the square root of three squared minus two squared is the square root of nine minus four is the square root of five

OpenStudy (anonymous):

@satellite73 can you write that in a equation please. I'm a visual learner not from reading words

OpenStudy (anonymous):

i wrote it in words because i was trying to convince you that you could do it by thinking aloud

OpenStudy (anonymous):

yeah I'm not that type of learner

OpenStudy (anonymous):

you can write \[\sqrt{3^2-2^2}=\sqrt{9-4}=\sqrt5\]

OpenStudy (anonymous):

okay then \[1=\sqrt{5}\]

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

@satellite73